The perimeter of a rectangle is 94 feet, and the area of the rectangle is 90ft^2. Find the dimensions of the rectangle.
Answer by jim_thompson5910(35256)

Perimeter

P = 2L+2W

94 = 2L+2W

94 - 2L = 2W

2W = -2L + 94

W = -2L/2 + 94/2

W = -L + 47

-------------------------------------------------------
Area

A = LW

A = L(-L + 47)

A = -L^2 + 47L

90 = -L^2 + 47L

0 = -L^2 + 47L - 90

-L^2 + 47L - 90 = 0

L^2 - 47L + 90 = 0

From here, solve the equation for L (using the quadratic formula). Make sure to toss out any negative answers.

Once you get a solution for L, use it to find W.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------